### v2

@@ 3.6 Interpretability of NEOS
$x_{t}=\sum_{i=1}^{N_{t}}x_{t,i}=\sum_{i=1}^{N_{t}}\sum_{j=1}^{T}p_{t,i,j}x_{t,i}=\sum_{j=1}^{T}\sum_{i=1}^{N_{t}}p_{t,i,j}x_{t,i}=\sum_{j=1}^{T}c_{t,j},$ (1)
where $x_{t}$ is NEOS in period $t$, $x_{t,i}$ is the score of article $i$ in period $t$, $N_{t}$ the number of articles in the respective period, $T$ the number of topics, $p_{t,i,j}$ the probability of an article $i$ to belong to topic $j$ in period $t$ (with $\sum_{j=1}^{T}p_{t,i,j}=1$) and $c_{t,j}$ the contribution of topic $j$ to $x_{t}$ in period $t$. Note that when decomposing the monthly standardized version of NEOS, the same mean and variance of the monthly indicator should be used to center and scale $x_{t,i}$ per article.
@@
A keyword-based method is useful to investigate the influence of specific ad-hoc topics, e.g., to quickly validate a policymaker’s intuition. By defining one or more keywords that must be present in an article for it to be considered relevant to a particular topic, we obtain $p_{t,i,j}~\epsilon~\{0,1\}$. In this simple form, this only allows us to distinguish between articles that belong to the defined topic and those that do not. Matching keywords is a crude approach to identify topics–articles which contain only one relevant section or where the entire article is relevant are weighed equally. Additionally, keyword-based methods lack context-awareness which can lead to incorrect interpretations.
@@
Classification-based methods are useful for pre-defined topics. A common approach is to label a subset of the data by humans or LLMs for training (e.g. bucur:2025). To define the topics, expert knowledge is required (see Section 3.3 for a discussion of the advantages and disadvantages of manual labeling). The topic labels are then used to train a classifier, using the embeddings as input. We exploit the fact that we instructed the LLM to generate synthetic articles for different economic sectors (labor market, trade, financial markets, see Section 3.3). Thus, our synthetic articles are already categorized by economic sectors. These categories can serve as topics. We train a regularized multinomial logistic regression on the embeddings of the synthetic example articles and apply the trained model to the embeddings of the news articles to obtain $p_{t,i,j}$. To simplify and ease interpretability, we limit the number of topics assigned to each article: We limit the total number of topics to a maximum of three, and select only the topics with the highest probabilities for which the cumulative sum of their probabilities in decreasing order exceeds 0.7. We rescale the probabilities of the selected topics per article such that they sum up to one again.
@@
Both of the above approaches rely on pre-defined topics. In contrast, automatically generated topics allow monitoring emerging trends in the data. To achieve this, we first apply UMAP (mcinnes:2020) to the embeddings of the news articles to get a new set of embeddings with lower dimensionality. Thereafter, we fit K-means clustering to the lower-dimensionality embeddings of the latest available month, i.e., November 2025. We then apply the resulting model to the historical data. This allows us to capture the latest developments while ensuring a consistent history. We choose K-means over more sophisticated clustering algorithms as the algorithm is deterministic (for a fixed seed), scalable and able to predict the cluster membership for new data points consistently.
@@ 4 Results
Figure 5 shows NEOS alongside with Swiss real year-on-year GDP growth and two traditional survey-based leading indicators (KOF Economic Barometer, Swiss Economic Confidence). NEOS closely tracks GDP growth and captures turning points similarly to established indicators. To assess the leading properties of NEOS, Appendix A.2 reports the correlations between NEOS and GDP growth at different lags alongside the same correlations for other traditional survey indicators for comparison. The correlations for NEOS are uniformly high and regularly exceed those of the survey indicators.
@@ 4.1 Forecast evaluation
Following the approach of kalamara:2022, we now evaluate the ability of NEOS and other survey-based indicators to improve forecasts of Swiss real GDP growth in a pseudo-out-of-sample evaluation over the period from Q1 1999 to Q1 2025.333Since recent releases of GDP are often subject to backward revisions, we limit the analysis to Q1 2025. The set-up is designed to mimic the information set of a policymaker: Towards the end of quarter $t$, the policymaker wants to forecast year-on-year GDP growth for the current quarter $t$ and for the next two quarters $t+1$ and $t+2$, while official GDP data are available only up to t1t-1.
For each horizon $h\in\{0,1,2\}$, we estimate the regression model
@@
where $y_{t}$ denotes year-on-year GDP growth in quarter $t$, $x_{t}^{(m)}$ is the value of a monthly indicator in month $m$ of quarter $t$, and $\varepsilon_{t}$ is a zero-mean error term.
We compare NEOS with several widely used monthly survey-based indicators: the manufacturing and service Purchasing Managers’ Index (PMI) for Switzerland, the KOF Business Situation Indicator, and the SECO Consumer Sentiment Index. Finally, we also consider a lexicon-based indicator constructed from the same news articles as NEOS. The lexicon-based indicator serves as a natural benchmark to compare against NEOS since it is based on exactly the same underlying data as NEOS but computed using a simpler approach (see Appendix A.3 for details).
All indicators are released at the end of each month or shortly thereafter. To reflect the information available to a policymaker towards the end of quarter $t$, we use as $x_{t}^{(m)}$ the value for the second month ($m=2$) of the quarter of NEOS or one of the competing indicators. In other words, forecasts for $y_{t+h}$ are based on GDP data up to t1t-1 and indicator information up to the second month of quarter $t$.
In addition, we also consider timelier versions of NEOS that rely on articles published within the first 7, 14, or 21 days of each month (see Section 3.5 for computation). As the monthly value of these NEOS versions can be computed before the end of the corresponding month, we can use it for the third month ($m=3$) of quarter $t$ in the forecasting exercise (in contrast to the other indicators, where we can only use the value for the second month as the value of the third month only becomes available in the “fourth” month, which is already the quarter ahead). This allows us to evaluate the additional forecast accuracy that can be gained by the possibility of observing NEOS in real time, well before most traditional data releases become available.
To assess the forecast performance for each indicator, we compare the regression model in Equation (2) to an autoregressive benchmark that only includes lagged GDP, i.e., Equation (2) with $\gamma=0$. This corresponds to an AR(1) model for year-on-year GDP growth. For each indicator and forecast horizon, we compute the root mean squared error (RMSE) of the pseudo-out-of-sample forecasts and report the ratio
@@
A value below one indicates that including the indicator as a predictor improves the accuracy of the forecast relative to the AR(1) model. We re-estimate the models recursively with an expanding window procedure starting with the first eight quarters of data. This means that each forecast is based only on information that would have been available at the time it was made.
We assess the statistical significance of the improvements in forecast accuracy using a modified Diebold–Mariano test. [^4] We follow beck:2025 and implement a modified version of the diebold:1995 test, using heteroskedasticity- and autocorrelation-consistent (HAC) standard errors to account for serial correlation in the forecast errors. Table 1 reports RMSE ratios, with Diebold–Mariano p-values in parentheses. The lowest RMSE ratio for each horizon is printed in bold.
@@
NEOS 0.864 (0.054) 0.802 (0.055) 0.818 (0.077)
NEOS (first 7 days in third month) 0.911 (0.067) 0.835 (0.069) 0.837 (0.085)
NEOS (first 14 days in third month) 0.910 (0.061) 0.797 (0.040) 0.812 (0.066)
NEOS (first 21 days in third month) 0.920 (0.049) 0.780 (0.032) 0.803 (0.072)
Lexicon-based approach 0.925 (0.148) 0.911 (0.164) 0.919 (0.218)
Manufacturing PMI for Switzerland 0.826 (0.051) 0.800 (0.063) 0.817 (0.042)
Service PMI for Switzerlanda 0.851 (0.119) 0.963 (0.127) 0.953 (0.124)
KOF Business Situation Indicatorb 0.996 (0.452) 1.017 (0.904) 1.017 (0.828)
SECO Consumer Sentiment Indexc 0.965 (0.387) 0.989 (0.439) 0.969 (0.362)
RMSE ratios with Diebold–Mariano p-values in parentheses. Best RMSE values in bold.
a Evaluation from Q1 2014 to Q1 2025 due to availability.
b Evaluation from Q2 2009 to Q1 2025 due to availability.
c Quarterly indicator: GDP in $t+h$ regressed on indicator for quarter $t$.
NEOS and its timelier versions consistently improve forecasts across all horizons: the RMSE ratios are substantially below one, and the Diebold–Mariano p-values indicate that these gains are statistically significant at the 10% level. For the baseline version of NEOS, the RMSE ratios range between 0.80 and 0.86 across horizons, implying reductions in forecast errors of around 14–20% relative to the AR(1) benchmark. The timelier NEOS versions perform almost as well and sometimes even slightly better than standard monthly NEOS. This shows that much of the improvement in forecast accuracy can already be achieved several weeks before official data releases.
Compared to other indicators, NEOS performs competitively. The lexicon-based approach performs worse, with RMSE ratios close to one. The better performance of NEOS implies that the use of text embeddings and synthetic articles extract richer and more accurate macroeconomic information than basic lexicon-based methods. Among the traditional indicators, the manufacturing PMI is the only one with similar or slightly better performance than NEOS. However, PMI data are much costlier to collect and are released on a fixed schedule, whereas NEOS can be updated at essentially any frequency. The KOF Business Situation Indicator and the SECO Consumer Sentiment Index provide only limited improvements, with RMSE ratios around or above one for several horizons. Overall, these results indicate that NEOS provides an attractive combination of predictive performance and timeliness.
@@
To further highlight the value of NEOS during crises, Appendix A.4 shows that including NEOS in forecasts for GDP growth markedly reduces forecast errors during the global financial crisis and the COVID-19 pandemic.
We also find that our results are driven in part by the diversity of the media landscape. In our data, we observe the type of media outlet. We computed subindices of NEOS using publications from the financial press, national newspapers, regional newspapers or tabloid press only, or combinations of those and then evaluated the performance of these subindices. Interestingly, including all types of media seems to enhance the signal for forecasting GDP growth. Including the tabloid press for example generates more distinct peaks and troughs. This indicates that using a diverse variety of publications is preferable compared to using single outlets.

### v3

@@ 3.6 Interpretability of NEOS
$x_{t}=\sum_{i=1}^{N_{t}}x_{t,i}=\sum_{i=1}^{N_{t}}\sum_{j=1}^{T}p_{t,i,j}x_{t,i}=\sum_{j=1}^{T}\sum_{i=1}^{N_{t}}p_{t,i,j}x_{t,i}=\sum_{j=1}^{T}c_{t,j},$
where $x_{t}$ is NEOS in period $t$, $x_{t,i}$ is the score of article $i$ in period $t$, $N_{t}$ is the number of articles in the respective period, $T$ is the number of topics, $p_{t,i,j}$ is the probability than article $i$ belongs to topic $j$ in period $t$ (with $\sum_{j=1}^{T}p_{t,i,j}=1$) and $c_{t,j}$ is the contribution of topic $j$ to $x_{t}$ in period $t$. Note that when decomposing the monthly standardized version of NEOS, the same mean and variance of the monthly indicator should be used to center and scale $x_{t,i}$ per article.
@@
A keyword-based method is useful to investigate the influence of specific ad-hoc topics, e.g., to quickly validate a policymaker’s intuition. By defining one or more keywords that must be present in an article for it to be considered relevant to a particular topic, we obtain $p_{t,i,j}\in\{0,1\}$. In this simple form, this only allows us to distinguish between articles that belong to the defined topic and those that do not. Matching keywords is a crude approach to identify topics–articles which contain only one relevant section or where the entire article is relevant are weighed equally. Additionally, keyword-based methods lack context awareness which can lead to incorrect interpretations.
@@
Classification-based methods are useful for predefined topics. A common approach is to label a subset of the data by humans or LLMs for training (e.g. Bucur and Hill, 2025). To define the topics, expert knowledge is required (see Section 3.3 for a discussion of the advantages and disadvantages of manual labeling). The topic labels are then used to train a classifier, using the embeddings as input. We exploit the fact that we instructed the LLM to generate synthetic articles for different economic sectors (labor market, trade and financial markets; see Section 3.3). Thus, our synthetic articles are already categorized by economic sectors. These categories can serve as topics. We train a regularized multinomial logistic regression on the embeddings of the synthetic example articles and apply the trained model to the embeddings of the news articles to obtain $p_{t,i,j}$. To simplify and ease interpretability, we limit the number of topics assigned to each article: We limit the total number of topics to a maximum of three and select only the topics with the highest probabilities for which the cumulative sum of their probabilities in decreasing order exceeds 0.7. We rescale the probabilities of the selected topics per article such that they sum up to one again.
@@
Both of the above approaches rely on pre-defined topics. In contrast, automatically generated topics allow monitoring emerging trends in the data. To achieve this, we first apply UMAP (McInnes et al., 2020) to the embeddings of the news articles to obtain a new set of embeddings with lower dimensionality. Thereafter, we fit K-means clustering to the lower-dimensionality embeddings of the latest available month, i.e., November 2025. We then apply the resulting model to the historical data. This allows us to capture the latest developments while ensuring a consistent history. We select K-means over more sophisticated clustering algorithms as the algorithm is deterministic (for a fixed seed), scalable and able to predict the cluster membership for new data points consistently.
@@ 4 Results
Figure 5 shows monthly NEOS alongside real quarterly year-on-year GDP growth and two traditional, monthly, survey-based leading indicators (KOF Business Situation Indicator and Manufacturing Purchasing Managers’ Index). NEOS closely tracks GDP growth and captures turning points similarly to the established indicators. To demonstrate NEOS’s leading properties, we analyze the correlations between NEOS and GDP growth at different lags. Appendix  A.2 shows the correlations alongside the same correlations for other indicators, which are based on surveys or news text, for comparison. The correlations for NEOS are consistently high, often surpassing those of the other indicators. Next, we assess whether this strong co-movement of NEOS with GDP growth translates into improved forecasting performance.
In the remainder of this section, we conduct a comprehensive forecast evaluation of NEOS. We first compare its out-of-sample forecasting performance against a wide range of alternative indicators. We then examine the importance of NEOS’s timeliness by assessing how its predictive accuracy varies with different publication lags. Next, we analyze NEOS’s performance over time, highlighting how timely information helps during periods of crisis and high volatility. We also investigate our indicator’s incremental predictive power by testing whether NEOS adds significant information beyond established indicators in multivariate forecasting models.
In addition, we demonstrate how NEOS can be interpreted through an analysis of the underlying news content and sentiment drivers.
@@ 4.1 Forecast evaluation
Following the approach of Kalamara et al. (2022), we evaluate the ability of NEOS and other survey- or text-based indicators to improve real GDP growth forecasts in a pseudo-out-of-sample evaluation over the period from Q1 1999 to Q1 2025.333Since recent releases of GDP are often subject to backward revisions, we limit the analysis to Q1 2025. The set-up is designed to mimic the information set of a policymaker: Towards the end of quarter $t$, the policymaker wants to forecast year-on-year GDP growth for the current quarter $t$ and for the next two quarters $t+1$ and $t+2$, while official GDP data are available only up to t1t-1.
For each horizon $h\in\{0,1,2\}$, we estimate the following regression model
@@
where $y_{t}$ denotes the year-on-year GDP growth in quarter $t$, $x_{t}^{(m)}$ is the average of the indicator values from the first $m$ months of quarter $t$, and $\varepsilon_{t}$ is a zero-mean error term. We compare our indicator against a range of established monthly indicators. These include widely used survey-based measures: the manufacturing and services Purchasing Managers’ Indices (PMIs) for Switzerland, the KOF Business Situation Indicator, and the SECO Consumer Sentiment Index.
In addition, we evaluate NEOS against two alternative text-based indicators derived from the same underlying news data. Both alternative indicators use a lexicon/dictionary-based approach. The first is the well-established Economic Policy Uncertainty (EPU) index proposed by Baker et al. (2016). The second follows the methodology of Barbaglia et al. (2025) and relies on a German translation of their sophisticated lexicon. Both of these text-based indicators serve as a natural benchmark: they rely on exactly the same underlying news data but employ a more straightforward counting approach (see Appendix A.3 for details).
All survey indicators are released at the end of each month or shortly thereafter. To mimic the information set available to a policymaker near the end of quarter $t$, we define $x_{t}^{(m)}$ as the average of the indicator values from the first two months ($m=2$) of quarter $t$. Thus, forecasts of GDP growth $y_{t+h}$ rely on GDP observations up to quarter t1t-1 and indicator data available through the second month of quarter $t$.
A key advantage of NEOS is that it can be continuously updated as new articles become available. To evaluate the benefits of this feature, we construct variants that include articles published in the first 7, 14, or 21 days of the third month of quarter $t$ (see Section 3.5 for details). These variants incorporate more recent information and can be computed well before the end of the month—substantially earlier than conventional indicators. For these timelier variants of NEOS, we therefore set $m=3$, i.e. we use the first two months and additional indicator information available through the third month of quarter $t$ in the forecasting exercise. This allows us to quantify the improvement in the forecast accuracy arising from NEOS’s ability to exploit intra-month news flow and provide updates long before most traditional indicators are released.
To assess the forecast performance for each indicator, we compare the regression model in Equation (1) to an autoregressive benchmark that only includes lagged GDP, i.e., Equation (1) with $\gamma=0$. This corresponds to an AR(1) model for year-on-year GDP growth. For each indicator and forecast horizon, we compute the mean absolute error (MAE) of the pseudo-out-of-sample forecasts and report the ratio
@@
A value below one indicates that including the indicator as a predictor improves the accuracy of the forecast relative to the AR(1) model. We re-estimate the models using an expanding window, starting with an initial window containing the first eight quarters of data. Each forecast is based only on information that would have been available at the time it was made.
We assess the statistical significance of the improvements in forecast accuracy using a modified Diebold–Mariano test. [^4] We follow Beck and Wolf (2026) and implement a modified version of the Diebold and Mariano (1995) test, using heteroskedasticity- and autocorrelation-consistent (HAC) standard errors to account for serial correlation in the forecast errors. Table 1 reports the MAE ratios, with the Diebold–Mariano p-values in parentheses. The lowest MAE ratio for each horizon is printed in bold.
@@
NEOS 0.88 (0.03) 0.78 (0.06) 0.80 (0.05)
NEOS (first 7 days in the third month) 0.87 (0.03) 0.79 (0.06) 0.80 (0.05)
NEOS (first 14 days in the third month) 0.87 (0.03) 0.78 (0.06) 0.79 (0.04)
NEOS (first 21 days in the third month) 0.88 (0.03) 0.78 (0.06) 0.77 (0.03)
EPU for Switzerland 1.12 (0.09) 1.11 (0.89) 1.01 (0.52)
Lexicon-based approach 1.00 (0.49) 0.97 (0.41) 0.93 (0.30)
KOF Business Situation Indicator 1.06 (0.86) 1.01 (0.67) 1.00 (0.48)
SECO Consumer Sentiment Index 1.09 (0.80) 1.01 (0.53) 0.99 (0.47)
Manufacturing PMI for Switzerland 0.94 (0.22) 0.85 (0.16) 0.84 (0.09)
Services PMI for Switzerland 0.89 (0.10) 0.95 (0.10) 0.92 (0.12)
Notes: MAE ratios with the Diebold–Mariano p-values in parentheses. The best MAE values in bold. Due to limited data availability, the evaluation for the KOF Business Situation Indicator is limited to Q1 2009–Q1 2025 and that for the Services PMI is limited to Q1 2014–Q1 2025. In Appendix A.4, Tables A6 and A7, we provide the results for the subperiods Q1 2009–Q1 2025 and Q1 2014–Q1 2025. Since the SECO Consumer Sentiment Index is a quarterly indicator, we regress GDP in $t+h$ on the indicator for quarter $t$. Data sources: KOF Swiss Economic Institute, procure.ch, SECO, SNB, Swissdox@LiRI, UBS.
NEOS and its intra-month updated versions substantially improve forecast accuracy across all horizons, with MAE ratios ranging from 0.77 to 0.88 (implying error reductions of 12–23% relative to the AR(1) benchmark). These gains are statistically significant at the 10% level or better, as indicated by p-values below 0.06 in all cases.
The baseline NEOS achieves MAE ratios of 0.88 ($h=0$), 0.78 ($h=1$), and 0.80 ($h=2$). Incorporating news from the early days of the third month yields further improvements at longer horizons: the version using the first 14 days records 0.78 at $h=1$ and 0.79 at $h=2$, whereas the version using the first 21 days delivers the best performance at $h=1$ (0.78) and $h=2$ (0.77). At the nowcast horizon ($h=0$), the 7-day version is marginally superior (0.87). Overall, the intra-month variants match or outperform the baseline NEOS, demonstrating that substantial forecasting gains can be realized several weeks ahead of conventional indicator releases.
Relative to alternative indicators, NEOS dominates. The Economic Policy Uncertainty (EPU) index consistently yields MAE ratios above one. The alternative lexicon-based approach also yields MAE ratios close to one (0.93–1.00), with no statistically significant improvements. This highlights the value of NEOS’s embedding-based methodology and synthetic labeling procedure, which extracts substantially richer macroeconomic signals from the same underlying news articles.
The KOF Business Situation Indicator and SECO Consumer Sentiment Index also generally perform worse than the AR(1) benchmark (MAE ratios around or above one) and show no significant forecasting power. The Manufacturing and Services PMIs deliver moderate improvements (MAE ratios of 0.84-0.95) but perform worse than NEOS across all horizons. Notably, PMI surveys are resource-intensive and follow fixed release schedules, whereas NEOS leverages readily-available news data and can be updated continuously.
Due to limited data availability, the evaluation for the KOF Business Situation Indicator is limited to Q1 2009–Q1 2025 and for the Services PMI to Q1 2014–Q1 2025. In Appendix A.4, Tables A6 and A7, we provide the results for the respective subperiods. Overall, our findings carry over to these subsamples.
Taken together, the results establish that our methodology is highly competitive, combining strong predictive performance with exceptional timeliness.
4.2 Importance of timeliness
@@
4.3 Performance over time
Table 1 shows that including NEOS in the regression model in Equation (1) improves the accuracy of GDP growth forecasts relative to the AR(1) benchmark. To determine whether these gains are persistent or driven by outliers, we examine the path of the cumulative absolute error difference (CAED).
For a given horizon $h$, the CAED at time $T$ is defined as:
[…]
where $|e_{t,h}^{\mathrm{NEOS}}|$ and $|e_{t,h}^{\mathrm{AR}}|$ are the absolute forecast errors of the NEOS-augmented regression and the AR(1) benchmark, respectively. The CAED is computed over the same pseudo out-of-sample evaluation period used in Section 4.1. Under this specification, a downward-sloping trajectory indicates that the NEOS-augmented model outperforms the benchmark. In an ideal setting, the CAED series should be strictly downward sloping and remain below the zero threshold, indicating that the augmented model consistently incurs smaller forecast errors than the benchmark in every period.
Figure 7 displays the CAED for horizon $h=2$. The series declines most sharply during periods of macroeconomic distress, such as the Great Financial Crisis and the COVID-19 pandemic. Outside of these crisis episodes, the series continues to trend downward overall but features several extended flat segments. These flat periods predominantly coincide with intervals of relatively stable and constant GDP growth rates, during which the simple AR(1) benchmark is particularly difficult to outperform. In such low-volatility environments, the NEOS-augmented model delivers predictive accuracy roughly on par with the AR(1), yielding no material additional gains. Nevertheless, the model still provides meaningful predictive improvements during more turbulent periods, and the overall pattern of outperformance remains visible across the full sample. While Figure 7 focuses on $h=2$ the results for horizons $h=0$ and $h=1$ are provided in Appendix A.5; the qualitative patterns remain highly consistent across all three horizons.
Figure 7: Cumulative forecast improvements compared to an AR(1) model for forecast horizon $h=2$.
4.4 Kitchen sink regression
To assess the incremental predictive power of the NEOS indicator, we perform a ”kitchen sink” regression, including all established benchmarks in the model simultaneously. This test determines whether NEOS captures unique economic signals that are not already present in a linear combination of standard survey- and text-based measures. The regression equation is specified as follows:
$y_{t+h}=\alpha+\gamma\text{NEOS}_{t}+\beta_{1}y_{t-1}+\beta_{2}\text{mPMI}_{t}+\beta_{3}\text{SECO}_{t}+\beta_{4}\text{LEX}_{t}+\beta_{5}\text{EPU}_{t}+\epsilon_{t}.$
The control variables represent those indicators available over the full sample period and introduced in Section 4.1: the manufacturing Purchasing Managers’ Index (mPMI), the SECO Consumer Sentiment Index (SECO), and two text-based benchmarks derived from the same data source—the Economic Policy Uncertainty index (EPU) and a German-translated lexicon-based measure (LEX) following Barbaglia et al. (2025). The primary objective is to test the null hypothesis H_0: γ= 0. If we reject this hypothesis, it implies that the NEOS indicator provides statistically significant predictive value beyond the information contained in traditional surveys and existing dictionary-based text methods.
The inclusion of the COVID-19 period presents a significant challenge for stable parameter estimation. The extreme fluctuations in GDP growth during 2020 and 2021 act as massive outliers that can disproportionately influence the regression coefficients, potentially masking the underlying structural relationship between the indicators and economic activity. Following Schorfheide and Song (2024), who argued that excluding crisis observations is a robust and transparent alternative to complex outlier modeling in VAR contexts, we present the results both with and without the pandemic period to ensure the reliability of our estimates.
The results, presented in Table 2, broadly correspond to the earlier univariate findings but offer deeper insight into the indicator’s lead-lag structure. At the immediate horizon ($h=0$), the performance gap between NEOS and established indicators, specifically the manufacturing PMI, is relatively small, though the coefficient remains statistically significant when the COVID-19 period is excluded.
However, as the forecast horizon increases to $h=1$ and $h=2$, the relative strength of the NEOS indicator becomes more pronounced. Both the magnitude of the coefficient $\hat{\gamma}$ and its statistical significance increase at these longer horizons. This suggests that NEOS is particularly effective as a forward-looking measure; while survey indicators may reflect the current ”nowcast” environment, NEOS appears to capture underlying economic sentiment that anticipates shifts in GDP growth several quarters ahead. These results remain robust even when accounting for the high-volatility COVID-19 period, confirming that NEOS is a valuable and distinct addition to the Swiss economic forecasting toolkit.
Table 2: Incremental predictive power of the NEOS: kitchen sink regression results
$h=0$ $h=1$ $h=2$
$\hat{\gamma} -$0.001 0.002* 0.004* 0.005*** 0.006*** 0.005***
(0.002) (0.001) (0.002) (0.001) (0.002) (0.001)
Other Indicators
Include COVID-19
  • Notes: This table reports the estimated coefficient $\hat{\gamma}$ for the NEOS indicator in the kitchen sink regression. The model controls for lagged GDP growth, mPMI, SECO consumer sentiment, and text-based benchmarks (EPU and LEX). HAC standard errors are reported in parentheses. The results are presented including/excluding the COVID-19 period (Q1 2020 until Q4 2021). ***$p<0.01$, **$p<0.05$, *$p<0.1$.